the greater of two consecutive integers is 10 less than twice the smaller?
find the integers.

Answer by Paul(988)

Two consecutive intergers = x and x+1

(x+1)-10=2(x)
x-9=2x
-9=3x
x=-3

-3+1=-2

Hence the two integers are -2 and -3.
Paul.